A closed box, whose lenght is twice its width, is to have a surface area of 192 square inches. Find the dimensions of the box when the volume is maximum


1
Expert's answer
2021-01-11T15:03:32-0500


According to the question, the closed box represents a cuboid, which has a top as it is referred to as a closed box.


Let width of the box be "x" inches. Then the length of the box is 2"x" inches.


Now, using the formulae for the total surface area("S") of a cuboid we get:-



"S=2(lb+bh+lh)\\
\Rightarrow 2(2x.x+x.h+2x.h)=192\\
\Rightarrow 2(2x^2+xh+2xh)=192\\
\Rightarrow 2x^2+3xh=96\\
\Rightarrow 3xh=96-2x^2\\
\Rightarrow h=\frac{96}{3x}-\frac{2x^2}{3x}\\~\\
\Rightarrow h=\frac{32}{x}-\frac{2x}{3}"

So we get the height of the box, in terms of "x" .



Now, we have to have the volume("V") to be maximum.



"V=lbh\\~\\
\Rightarrow V=2x.x.(\frac{32}{x}-\frac{2x}{3})\\~\\
\Rightarrow V=2x^2(\frac{32}{x}-\frac{2x}{3})\\~\\
\Rightarrow V=64x-\frac{4x^3}{3}\\~\\"

Now, after differentiating both sides wrt x, we get:-



"\frac{dV}{dx}=64-4x^2\\~\\"

Again differentiating both sides wrt x, we get:-



"\frac{d^2V}{dx^2}=-8x"

Now, as we clearly see that, "\frac{d^2V}{dx^2}" is always negative (as x is a dimension and cannot be negative), so "\frac{dV}{dx}=0" , will give us a value of x which makes V maximum.


Now,



"\frac{dV}{dx}=0\\
\Rightarrow 64-4x^2=0\\
\Rightarrow 4x^2=64\\
\Rightarrow x^2=16\\
\Rightarrow x=\pm4"

Here, we neglect the value "x=-4" as we know, that the width of a box cannot be negative.


So, we have "x=4" , when the volume of the box is maximum.


Therefore the dimensions of the box are:-


length = "2x=2(4)=8" inches.

width = "x=4" inches.

height = "\frac{32}{x}-\frac{2x}{3}=\frac{32}{4}-\frac{2(4)}{3}=8-\frac{8}{3}=\frac{16}{3}" inches.


and, the maximum volume(V) of the box is:- "64(4)-\frac{4(4)^3}{3}=256-\frac{256}{3}=\frac{512}{3}" cubic inches.